OpenStudy (anonymous):

Given the function f(x)=2/x+1, evaluate f(-5)

OpenStudy (anonymous):

Put -5 instead of x It will be F(-5)=2/-5+1 ..and u evaluate it ..

OpenStudy (anonymous):

so -5=1

OpenStudy (anonymous):

F(-5)=2/-5+1 F(-5)=-2/5+1 F(-5)=3/5

OpenStudy (anonymous):

why does it keep a 5 as a denom.

OpenStudy (anonymous):

so it would be x=3/5